Find the equation of integral surface to the differential equation
2y(z-3)p+(2x-z)q=y(2x-3)
1
Expert's answer
2020-10-06T14:15:38-0400

"2y(z-3)p+(2x-z)q=y(2x-3)\\

\displaystyle p = \frac{\partial z}{\partial x}, q = \frac{\partial z}{\partial y}\\

\textsf{The equation for the total}\\\textsf{differential is thus given as}\\


\mathrm{d}z = \frac{\partial z}{\partial x} \mathrm{d}x + \frac{\partial z}{\partial y}\mathrm{d}y \\


\textsf{The differentials are;}\\
\begin{aligned}
\mathrm{d}z &= 2xy - 6y\\
\mathrm{d}x &= 2yz - 6y\\
\mathrm{d}y &= 2x - z
\end{aligned} \\

\textsf{The Lagrange’s auxiliary equations}\\\textsf{for given PDE is}\\

\frac{\mathrm{d}z}{y(2x - 3)}= 
\frac{\mathrm{d}y}{(2x - z)} = \frac{\mathrm{d}x}{2y(z-3)} \\


\textsf{Choosing}\hspace{0.1cm} (x,3y,−z) \hspace{0.1cm}\\ \textsf{as multipliers, we have}\\

x\mathrm{d}x + 3y\mathrm{d}y - z\mathrm{d}z = 0\\

\textsf{Integrating both sides, we have thus,}\\

x² + 3y² - z² = C_1 \\

\textbf{\textsf{Or}}\\

\textsf{Choosing}\hspace{0.1cm} (l,3m,−n) \hspace{0.1cm} \\\textsf{as multipliers, we have}\\


l\mathrm{d}x + 3m\mathrm{d}y - n\mathrm{d}z = 0\\

\Rightarrow lx + 3my - nz = C_2.\\

\textsf{Thus, the solution of the given PDE is}\\
\phi(lx + 3my - nz, x² + 3y² - z²) = 0 \hspace{0.2cm}\forall l,m,n \in \mathbb{R}\hspace{0.1cm}\&\\

C_1 \hspace{0.1cm}\textsf{and}\hspace{0.1cm}C_2\hspace{0.1cm}\textsf{are arbitrary constants}"